### v2

@@ 3.3 Ablation Study
For the remaining evaluation metrics—Accuracy (ACC), Balanced Accuracy (BACC), Sensitivity (SEN), Specificity (SPEC), and Matthews Correlation Coefficient (MCC)—we highlight the top two performing models for each metric. The best model is marked in red, and the second-best model in blue. The results of these three evaluation methods are shown in Figure 5.
Results and Analysis
1. ROC/PR Curves
The full model demonstrates strong performance across all tasks, with median AUC of 0.991 (80% of tasks achieving AUC between 0.982-0.998). Key observations:
- Removing sequence embeddings causes 9.7% AUC drop (from 0.985 to 0.890) in PTM predictions
- Structure embedding removal shows smaller impact (3.2% average AUC decrease)
- Dual component removal leads to maximum 12.7% performance reduction
2. Confusion Matrix
Component removal creates distinct error patterns:
- Missing CNN increases false alarms by 2.1× (e.g., antimicrobial detection: 84→176 cases),which suggests that the CNN module plays a critical role in filtering out incorrect predictions
- Losing Transformer raises missed detections by 37.5% (anticancer task: 152→209 cases),highlighting the importance of this module in correctly identifying positive instances.
- Positional encoding removal causes combined degradation (both error types increasing over 100%)
This behavior is consistent across different datasets, as indicated by the confusion matrix behavior for each ablation variant.
3. UMAP Visualization: As depicted in Figure 6& SuppFig 3, the UMAP projections clearly show that the full model achieves superior separation between positive and negative samples in the feature space. In contrast, removing embeddings or key modules results in overlapping distributions, demonstrating a reduction in the model’s ability to discriminate between the classes.
- Average distance between categories: 1.73 vs 0.91 in others
- Category differentiation improved by 29%
- Decision boundaries 3× clearer
These advantages enable 6.8% better accuracy on challenging cases.
4. Average Performance Metrics In this part of the analysis, we focus on the performance of different models across multiple evaluation metrics. By comparing the baseline model (PDeepPP) and the six ablation variants on average values for each metric, we can better understand the contribution of each key module.
First, Accuracy (ACC) and Balanced Accuracy (BACC) exhibit similar trends. From the table, we see that the baseline model (PDeepPP) achieves the highest scores for these two metrics, 0.8349 and 0.8282, respectively. Removing the embedding module results in almost the same performance as the baseline, indicating that embeddings have a minimal effect on these metrics. However, after removing the attention mechanism and translinear layer, both metrics show a noticeable decline, particularly BACC, with scores of 0.8267 and 0.8359, slightly lower than the baseline. This suggests that the attention mechanism and translinear layer play a significant role in improving model performance.
For Sensitivity (SN) and Specificity (SP), the baseline model and the embedding-removed model have identical scores, indicating that the embedding module has a relatively mild impact on these metrics. Removing the CNN module results in a reduction in sensitivity (0.7915), while specificity remains at a relatively high level (0.8301). This indicates that removing the CNN module increases false negatives (FN), affecting the model’s ability to recognize positive samples. In contrast, removing the Transformer module leads to an increase in false positives (FP), resulting in a higher specificity but a more significant decline in sensitivity.
Matthews Correlation Coefficient (MCC) shows a similar trend. The MCC value for the baseline model (PDeepPP) is 0.8265, the highest among all models. The models without embeddings and without CNN modules follow closely, with scores of 0.8265 and 0.8089, respectively. In contrast, removing the PosEncoding and PosCNN modules results in a substantial drop in MCC values, 0.6511 and 0.5503, respectively, indicating that removing these modules severely affects the model’s overall performance.
Overall, the baseline model (PDeepPP) performs best across all metrics, while removing the embedding module has the least effect on performance. On the other hand, removing PosEncoding and PosCNN leads to the most significant performance degradation. In particular, removing PosEncoding and PosCNN not only reduces accuracy but also significantly affects sensitivity and MCC, indicating that they play a critical role in the model.
4 Conclusion
The PDeepPP framework represents a significant leap forward in the field of peptide identification and PTM prediction, showcasing the transformative power of deep learning in addressing complex bioinformatics challenges. Our experimental results demonstrate that PDeepPP consistently outperforms state-of-the-art models across a variety of tasks, underscoring its robustness and adaptability. For instance, in antimicrobial peptide identification, PDeepPP achieved an impressive accuracy (ACC) of 0.9726 and a PR AUC of 0.9977, significantly surpassing the performance of UniDL4BioPep. Similarly, in phosphorylation site prediction, a critical PTM task, PDeepPP achieved an ACC of 0.9984 and a PR AUC of 0.9910, highlighting its exceptional predictive capabilities and ability to generalize across diverse datasets. Furthermore, PDeepPP substantially reduced false positives (FP) and false negatives (FN) across all tested datasets, reinforcing its practical utility in real-world applications.
PDeepPP’s success can be attributed to its innovative integration of pretrained protein language models (ESM-2) with a parallel neural network architecture combining Transformers and CNNs. This design allows the model to capture both local and global sequence features, enhancing its ability to handle the complexity of protein sequences. The inclusion of the Transductive Information Maximization (TIM) loss function further improves performance on imbalanced datasets, a common challenge in bioinformatics. This combination of advanced deep learning techniques enables PDeepPP to achieve superior accuracy, robustness, and scalability compared to traditional methods.
Compared to conventional models, PDeepPP significantly reduces the need for extensive feature engineering and manual annotation, making it adaptable to a wide range of PTM types and peptide functionalities. Its ability to handle large-scale datasets and diverse biological tasks positions it as a valuable tool for large-scale bioinformatics research. Moreover, PDeepPP’s computational efficiency and reduced reliance on experimental data make it a cost-effective solution for peptide discovery and PTM analysis.
Looking ahead, PDeepPP holds immense potential for broader applications in both research and industry. Future directions include integrating structural and spatial data to further enhance predictive accuracy, as well as combining the framework with high-throughput experimental workflows to accelerate large-scale protein function discovery and drug development. Additionally, as novel PTM types, such as AMPylation, are discovered, PDeepPP can provide computational support for exploring these uncharted areas, advancing its potential applications in precision medicine and therapeutic innovation. The framework could also be extended to address multi-label identification tasks, where peptides may exhibit multiple bioactivities or PTMs simultaneously, further broadening its utility.
In summary, PDeepPP combines the strengths of pretrained protein language models, advanced neural network architectures, and innovative loss functions to offer a cutting-edge solution for protein sequence analysis. It represents a major advancement in the field of bioinformatics and has the potential to play a pivotal role in accelerating peptide discovery, PTM functional analysis, and the development of novel therapeutic strategies. By reducing the reliance on costly and time-consuming experimental methods, PDeepPP not only enhances the efficiency of bioinformatics research but also opens new avenues for understanding the complex mechanisms underlying protein modifications and peptide bioactivity. As the field continues to evolve, PDeepPP is poised to drive innovation in life sciences, contributing to the development of targeted therapies and personalized medicine.
Data and Code Availability

### v3

@@ 8.4 Detailed Ablation Experimental Setup
For the remaining evaluation metrics—Accuracy (ACC), Balanced Accuracy (BACC), Sensitivity (SEN), Specificity (SPEC), and Matthews Correlation Coefficient (MCC)—we highlight the top two performing models for each metric. The best model is marked in red, and the second-best model in blue.
8.5 UMAP Analysis of Ablation Study
This analysis explores the contributions of PDeepPP’s key components to feature representation, utilizing Uniform Manifold Approximation and Projection (UMAP) visualization. By observing changes in the 2D feature space distribution across eight biological tasks, we can assess the synergistic effects of the model’s hybrid design. The analysis compares the baseline PDeepPP model against ablated versions, which are grouped into the following three categories:
  • 1.
    Embedding & Loss Function: The BaseEmbedding module or the TIM loss is removed.
  • 2.
    Transformer Module: The TransLinear module or its attention mechanism is removed.
  • 3.
    CNN Module: The PosCNN module or its positional encoding is removed.
An effective feature space is characterized by high intra-class compactness and clear inter-class separation.
8.5.1 Anticancer main
For this dataset, the baseline PDeepPP model exhibits the most defined clusters. Ablating any core component degrades performance. The removal of either the TransLinear module (Category 2) or the PosCNN module (Category 3) leads to a significant loss of class separation, with substantial mixing between positive and negative points, indicating that both global and local feature extraction paths are crucial. Similarly, removing the TIM loss function (Category 1) causes clusters to become diffuse and overlap considerably, highlighting its importance. The removal of the task-adaptive BaseEmbedding also degrades performance, though to a lesser extent than ablating the core modules or the loss function.
8.5.2 Antiviral
Consistent with the anticancer task, the baseline model provides the best class separation. Removing the TransLinear module (Category 2) or the PosCNN module (Category 3) again results in a marked decrease in cluster quality, underscoring the importance of both global context and local features for identifying antiviral peptides. The ablation of the TIM loss (Category 1) also leads to a significant increase in the overlapping region, while the removal of the embedding has a less pronounced, intermediate effect.
8.5.3 N6-acetyllysine
On this imbalanced PTM dataset, the baseline model forms two relatively clear clusters. The removal of the TIM loss function (Category 1) is particularly detrimental, causing the sample points to merge into a single, largely indistinguishable cluster. Likewise, ablating either the TransLinear module (Category 2) or the PosCNN module (Category 3) results in a substantial loss of feature space structure, suggesting that both global features and local conserved motifs are prerequisites for effective prediction.
8.5.4 Ubiquitination
The trend observed in previous tasks is consistent here, with the PDeepPP baseline providing the clearest boundaries. Removing the TIM loss function (Category 1) has a more pronounced negative impact on clustering quality than removing the embedding layer. The removal of either the Transformer path (Category 2) or the CNN path (Category 3) also leads to a significant degradation in the class structure, confirming both are important for learning effective representations.
8.5.5 methylation-G
For this PTM type, the baseline PDeepPP model provides the most effective clustering. Ablating either the TransLinear module (Category 2) or the PosCNN module (Category 3) leads to a noticeable degradation of the cluster structure, indicating the model benefits from both global sequence information and local features. Consistent with other tasks, removing the TIM loss function (Category 1) also significantly degrades class separation.
8.5.6 methylation-R
The PDeepPP baseline again shows the best performance on this dataset. The TIM loss function (Category 1) proves to be critical for achieving clear separation. The degradation of the feature space upon removing either the global feature path (TransLinear, Category 2) or the local feature path (PosCNN, Category 3) confirms that both global context and local motifs are valuable components for this prediction task.
8.5.7 Ubiquitin_K*
The baseline PDeepPP model demonstrates superior clustering for this common modification. A significant loss of separation is observed upon removing the TIM loss function (Category 1). Furthermore, ablating the TransLinear module (Category 2) or the PosCNN module (Category 3) leads to a substantial mixing of sample points, indicating that both global features and local motifs are key contributors to the model’s performance.
8.5.8 Crotonylation_K
For this less common modification, the baseline model’s clustering is well-separated. The TIM loss function (Category 1) is again shown to be effective, as its removal leads to poor clustering. The ablation of the TransLinear module (Category 2) confirms the importance of global dependencies, while the significant drop in cluster quality upon removing the PosCNN module (Category 3) indicates that local, conserved features are equally vital for identifying crotonylation sites.
Figure 1: The remaining ten tasks showing consistent performance in comparison with the UniDL4BioPep model. The upper section shows the AUC curves and confusion matrix combinations for each task. In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the UniDL4BioPep model. The color intensity of each block in the matrix indicates the proportion of that class relative to the total negativ e/positive samples.The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP. The label order is: (A) ACE, (B) DPPIV, (C) bitter, (D) Quorum, (E) Anticancer_alternative, (F) Anti-MRSA, (G) Antiparasitic, (H) neuro, (I) Toxicity, (J) Antioxidant.
Figure 2: The Remaining six tasks showing consistent performance in comparison with the MusiteDeep model. The upper section shows the AUC curves and confusion matrix combinations for each task. In the confusion matrices, the numbers in parentheses next to the values represent the difference in the count of this metric compared to PDeepPP for the MusiteDeep model. The color intensity of each block in the matrix indicates the proportion of that class relative to the total negative/positive samples. The lower section displays the UMAP plots for each dataset after feature extraction by PDeepPP, with the site distribution of these six datasets shown in the bottom right. The label order is: (A) Phosphoserine_Phosphothreonine_S,T, (B) N-linked-glycosylation_N, (C) O-linked-glycosylation_S,T, (D) Methylation_R, (E) S-Palmitoylation_C, (F) SUMOylation_K, (G) Hydroxyproline_P.
Figure 3: Four datasets selected from existing datasets, with respect to the given category (see legend). UMAP plots are shown for each pair of related ablated models and the baseline model (PDeepPP) after feature extraction. (A)-(D) correspond to methylation-G, methylation-R, Ubiquitin-K*, and Crotonylation-K, respectively.
Table 1: Benchmark Datasets Sourced from Publications Featuring State-of-the-Art Models
Tasks Training dataset Test dataset Reference
Bioactivity Positives Negatives Positives Negatives
ACE inhibitory activity 913 913 386 386 [39]
DPP IV inhibitory activity 532 532 133 133 [40]
Bitter 256 256 64 64 [41]
Umami 112 241 28 61 [42]
Antimicrobial activity 3876 9552 2584 6369 [43]
Antimalarial activity main 111 1708 28 427 [44]
alternaive 111 542 28 135 [37]
Quorum sensing activity 200 200 20 20 [45]
Anticancer activity main 689 689 172 172 [46,47]
alternative 776 776 194 194 [6]
Anti-MRSA strains activity 118 678 30 169 [48]
Tumor T cell antigens 470 318 122 75 [49]
Blood–Brain Barrier 100 100 19 19 [50]
Antiparasitic activity 255 255 46 46 [51]
Neuropeptide 1940 1940 485 485 [52,53]
Antibacterial activity 6583 6583 1695 1695 [54]
Antifungal activity 778 778 215 215 [54]
Antiviral activity 2321 2321 623 623 [54]
Toxicity 1642 1642 290 290 [55]
Antioxidant activity 582 541 146 135 [56]
PTMs[23] Positives Negatives Positives Negatives Reference
Phosphoserine/threonine 25170 473607 4847 96667
Phosphotyrosine 6939 64884 1669 17123
N-linked glycosylation 52926 318092 12836 76512
O-linked glycosylation 567 28446 143 6610
N6-acetyllysine 16222 199649 3895 48653
Methylarginine 3749 85090 966 20825
Methyllysine 1324 25836 335 7640
S-palmitoylation-cysteine 2260 11904 541 3172
Pyridoxine-carboxylic-acid 1128 7688 285 1554
Ubiquitination 2528 26772 581 7797
SUMOylation 795 16209 218 4036
Hydroxylysine 390 2968 121 661
Hydroxyproline 3931 13910 892 3631
Remaining PTMs Positives Negatives Positives Negatives Reference
methylation-G 627 10490 165 2615 [58]
methylation-R 1038 1038 290 290 [59]
Ubiquitin_K* 2528 26772 581 7797 [60]
Crotonylation_K 6975 6975 3989 3989 [61]
[…]
Table 2: Comparison with UniDL4BioPep in bioactive peptides and state-of-the-art models from the same benchmark datasets
Bioactivity Model name ACC ROC PR BACC Sn Sp MCC
ACE inhibitory activity PDeepPP 0.8225 0.9049 0.9144 0.8225 0.8782 0.7668 0.6491
UniDL4BioPep 0.8433 0.8918 0.8728 0.8433 0.8238 0.8627 0.6870
mAHTPred 0.883 0.951 N/A 0.894 0.873 0.767
DPP IV inhibitory activity PDeepPP 0.8947 0.9615 0.9654 0.8947 0.8496 0.9398 0.7927
UniDL4BioPep 0.8534 0.9384 0.9444 0.8534 0.8647 0.8421 0.7069
iDPPIV-SCM 0.797 0.847 N/A 0.789 0.805 0.594
Bitter PDeepPP 0.9141 0.9625 0.9692 0.9141 0.8906 0.9375 0.8290
UniDL4BioPep 0.9375 0.9824 0.9838 0.9375 0.9219 0.9531 0.8754
BERT4Bitter N/A 0.922 0.938 0.906 0.844 0.964
iBitter-Fuse 0.93 0.933 N/A 0.938 0.922 0.859
Umami PDeepPP 0.9213 0.9543 0.9835 0.9330 0.9016 0.9643 0.8325
UniDL4BioPep 0.8764 0.9417 0.9760 0.8326 0.9508 0.7143 0.7055
UniDL4BioPep-FL 0.888 0.943 0.883 0.892 0.875 0.733
iUmami-SCM 0.865 0.898 N/A 0.714 0.934 0.679
Antimicrobial activity PDeepPP 0.9726 0.9947 0.9977 0.9609 0.9887 0.9330 0.9329
UniDL4BioPep 0.9631 0.9910 0.9959 0.9532 0.9768 0.9296 0.9099
UniDL4BioPep-FL 0.96 0.991 0.961 0.96 0.963 0.903
TransImbAMP N/A N/A 0.969 0.963 0.974 N/A
Antimalarial activity (main dataset) PDeepPP 0.9758 0.9049 0.9898 0.8203 0.9977 0.6429 0.7695
UniDL4BioPep 0.9780 0.8898 0.9872 0.8214 1.0000 0.6429 0.7926
UniDL4BioPep-FL 0.978 0.898 0.965 0.979 0.95 0.793
iAMAP-SCM 0.978 0.82 0.826 0.654 0.998 0.776
Antimalarial activity (alternative dataset) PDeepPP 0.9877 0.9929 0.9984 0.9643 1.0000 0.9286 0.9566
UniDL4BioPep 0.9816 0.9844 0.9965 0.9464 1.0000 0.8929 0.9346
UniDL4BioPep-FL 0.989 0.987 0.993 0.985 1 0.9570
iAMAP-SCM 0.957 0.903 0.896 0.808 0.985 0.834
Quorum sensing activity PDeepPP 0.9750 0.9975 0.9976 0.9750 0.9500 1.0000 0.9512
UniDL4BioPep 0.9500 0.9900 0.9908 0.9500 0.9000 1.0000 0.9045
iQSP 0.93 0.96 N/A N/A N/A 0.86
QSPred-FL 0.943 0.945 N/A 0.935 0.95 0.885
Anticancer activity (main dataset) PDeepPP 0.7587 0.8334 0.8308 0.7587 0.8547 0.6628 0.5272
UniDL4BioPep 0.7355 0.8054 0.7780 0.7355 0.7326 0.7384 0.4709
iACP-FSCM 0.825 0.812 0.825 0.726 0.903 0.646
AntiCP2.0 0.754 N/A 0.754 0.774 0.734 0.51
Anticancer activity (alternative dataset) PDeepPP 0.9433 0.9709 0.9668 0.9433 0.9639 0.9227 0.8874
UniDL4BioPep 0.9459 0.9711 0.9630 0.9459 0.9794 0.9124 0.8938
iACP-FSCM 0.889 0.93 N/A 0.876 0.902 0.779
AntiCP2.0 0.92 N/A N/A 0.923 0.918 0.84
Anti-MRSA strains activity PDeepPP 0.9950 0.9998 1.0000 0.9833 1.0000 0.9667 0.9803
UniDL4BioPep 0.9899 0.9986 0.9998 0.9667 1.0000 0.9333 0.9604
UniDL4BioPep-FL 0.994 0.999 0.997 0.994 1 0.98
SCMRSA 0.96 0.986 0.935 0.9 0.97 0.848
Tumor T cell antigens PDeepPP 0.7563 0.8198 0.7662 0.7185 0.5600 0.8770 0.4680
UniDL4BioPep 0.7513 0.7883 0.7395 0.7144 0.5600 0.8689 0.4569
UniDL4BioPep-FL 0.746 0.796 0.762 0.734 0.791 0.446
iTTCA-Hybrid 0.71 0.756 N/A 0.844 0.493 0.363
Blood–Brain Barrier PDeepPP 0.9211 0.9640 0.9680 0.9211 0.9474 0.8947 0.8433
UniDL4BioPep 0.8421 0.9224 0.9211 0.8421 0.8947 0.7895 0.6880
BBBpred 0.7895 0.7895 N/A 0.6316 0.9474 0.6102
Antiparasitic activity PDeepPP 0.7609 0.9267 0.9370 0.7609 0.9565 0.5652 0.5669
UniDL4BioPep 0.8478 0.9045 0.9258 0.8478 0.8043 0.8913 0.6983
PredAPP 0.88 0.922 N/A 0.978 0.783 0.775
Neuropeptide PDeepPP 0.9021 0.9611 0.9602 0.9021 0.9010 0.9031 0.8041
UniDL4BioPep 0.9052 0.9707 0.9707 0.9052 0.8784 0.9320 0.8115
PreNeuroP 0.897 0.954 N/A 0.886 0.907 0.794
NeuroPred-CLQ 0.936 0.988 N/A 0.897 0.975 0.875
Antibacterial activity PDeepPP 0.9478 0.9811 0.9752 0.9478 0.9617 0.9339 0.8959
UniDL4BioPep 0.9395 0.9775 0.9704 0.9395 0.9687 0.9103 0.8806
ABPDiscover 0.935 0.975 N/A 0.912 0.957 0.87
Antifungal activity PDeepPP 0.9535 0.9911 0.9916 0.9535 0.9442 0.9628 0.9071
UniDL4BioPep 0.9512 0.9862 0.9838 0.9512 0.9628 0.9395 0.9026
ABPDiscover 0.942 0.988 N/A 0.921 0.963 0.884
Antiviral activity PDeepPP 0.8531 0.9192 0.9164 0.8531 0.9486 0.7576 0.7195
UniDL4BioPep 0.8291 0.9012 0.8827 0.8291 0.9037 0.7544 0.6656
ABPDiscover 0.828 0.896 N/A 0.764 0.892 0.662
Toxicity PDeepPP 0.9172 0.9779 0.9770 0.9168 0.9108 0.9228 0.8336
UniDL4BioPep 0.9603 0.9943 0.9934 0.9608 0.9665 0.9550 0.9205
ATSE 0.952 0.976 N/A 0.965 0.94 0.903
Antioxidant activity PDeepPP 0.7972 0.8776 0.8880 0.7981 0.7740 0.8222 0.5959
UniDL4BioPep 0.8221 0.9296 0.9410 0.8229 0.8014 0.8444 0.6454
AntiOxPred-FRS N/A 0.79 N/A N/A N/A 0.48
Table 3: Comparison with the MusiteDeep and PhosF3C in PTMs from the same benchmark datasets
PTM types Model ACC ROC PR BACC SN SP MCC
Hydroxyproline PDeepPP 0.9885 0.9992 0.9959 0.9729 0.9504 0.9955 0.9557
MusiteDeep 0.9412 0.9937 0.9718 0.9618 0.9917 0.9319 0.8187
PhosF3C 0.9476 0.9625 0.8892 0.8745 0.7686 0.9803 0.7912
Hydroxylysine PDeepPP 0.9648 0.9915 0.9615 0.9430 0.9070 0.9791 0.8887
MusiteDeep 0.8720 0.9866 0.9523 0.9152 0.9865 0.8438 0.7084
PhosF3C 0.8998 0.9473 0.8725 0.8645 0.8061 0.9229 0.6992
Methyllysine PDeepPP 0.9984 0.9997 0.9972 0.9924 0.9859 0.9990 0.9809
MusiteDeep 0.9877 0.9921 0.9649 0.9842 0.9803 0.9881 0.8756
PhosF3C 0.9869 0.9768 0.9005 0.9448 0.8986 0.9910 0.8527
Methylarginine PDeepPP 0.9918 0.9977 0.9565 0.9400 0.8830 0.9969 0.9016
MusiteDeep 0.9621 0.9949 0.9638 0.9767 0.9928 0.9607 0.7171
PhosF3C 0.9794 0.9702 0.8277 0.8969 0.8064 0.9875 0.7665
N-linked glycosylation PDeepPP 0.9810 0.9970 0.9806 0.9644 0.9415 0.9873 0.9204
MusiteDeep 0.9597 0.9883 0.8874 0.9760 0.9986 0.9535 0.8579
PhosF3C 0.9636 0.9869 0.8812 0.9740 0.9884 0.9596 0.8674
N6-acetyllysine PDeepPP 0.9870 0.9959 0.9589 0.9491 0.9045 0.9936 0.9049
MusiteDeep 0.9645 0.9939 0.9105 0.9760 0.9895 0.9625 0.8033
PhosF3C 0.9687 0.9574 0.8620 0.9041 0.8282 0.9800 0.7807
Phosphotyrosine PDeepPP 0.9944 0.9989 0.9909 0.9818 0.9664 0.9971 0.9654
MusiteDeep 0.9567 0.9943 0.9417 0.9711 0.9886 0.9536 0.7967
PhosF3C 0.9486 0.9542 0.8210 0.8964 0.8328 0.9599 0.7194
Phosphoserine/threonine PDeepPP 0.9647 0.9762 0.6487 0.7774 0.5702 0.9845 0.5896
MusiteDeep 0.8180 0.9386 0.4441 0.8749 0.9379 0.8120 0.3836
PhosF3C 0.9373 0.9841 0.7452 0.9522 0.9686 0.9357 0.6227
Pyrrolidone-carboxylic-acid PDeepPP 0.9804 0.9964 0.9822 0.9626 0.9368 0.9884 0.9253
MusiteDeep 0.9625 0.9978 0.9910 0.9735 0.9895 0.9575 0.8749
PhosF3C 0.9641 0.9772 0.9317 0.9358 0.8947 0.9768 0.8642
O-linked glycosylation PDeepPP 0.9947 0.9457 0.8231 0.8878 0.7762 0.9994 0.8631
MusiteDeep 0.9508 0.9916 0.9193 0.9680 0.9860 0.9501 0.5291
PhosF3C 0.9730 0.8720 0.5597 0.8186 0.6573 0.9799 0.5090
S-palmitoylation-cysteine PDeepPP 0.9908 0.9971 0.9906 0.9885 0.9852 0.9918 0.9639
MusiteDeep 0.9793 0.9965 0.9861 0.9764 0.9723 0.9805 0.9207
PhosF3C 0.9666 0.9871 0.9453 0.9544 0.9372 0.9716 0.8728
SUMOylation PDeepPP 0.9915 0.9982 0.9713 0.9521 0.9083 0.9960 0.9123
MusiteDeep 0.9838 0.9987 0.9755 0.9828 0.9817 0.9839 0.8600
PhosF3C 0.9831 0.9459 0.8725 0.9086 0.8257 0.9916 0.8245
Ubiquitination PDeepPP 0.9922 0.9983 0.9794 0.9656 0.9346 0.9965 0.9394
MusiteDeep 0.9383 0.9819 0.8969 0.9453 0.9535 0.9372 0.6851
PhosF3C 0.9511 0.9522 0.8397 0.8965 0.8330 0.9599 0.6866
[…]
Table 4: Model Hyperparameter Configuration
Category Parameter Value Description
Training Config Learning Rate 0.001 Initial learning rate for Adam optimizer
Batch Size 32/128 Training batch size (32) and inference batch size (128)
Training Epochs 100 Maximum training iterations with early stopping
Patience 20 Epochs tolerance for validation loss plateau
Architecture ESM-2 Dimension 1280 Output dimension of pretrained ESM-2 embeddings
Transformer Dim 128 Hidden dimension in self-attention layers
Attention Heads 8 Parallel attention mechanisms per layer
Encoder Layers 4 Stacked transformer encoder blocks
CNN Channels 64→32→64 Feature channel progression in convolutional modules
Regularization Dropout 0.3/0.15 Dropout rates: attention (0.3), CNN (0.15)
Position Encoding Yes Learnable positional encodings
Weight Decay None No explicit L2 regularization applied
UMAP Visualization n_neighbors 15 Local neighborhood size for manifold approximation
min_dist 0.1 Minimum embedding distance between points
random_state 42 Seed for reproducible visualization
Loss Function $\lambda$ [0.9,1.0] Cross-entropy weight in TIM loss
$\alpha$ 1.0 Entropy regularization coefficient
Threshold 0.5 Decision boundary for identification
Feature Fusion ESM Ratio [0.9,1.0] Weighting between ESM-2 and task-specific embeddings
  • Notes:
    1. 1.
      Final testing uses optimal parameters
    2. 2.
      UMAP parameters specifically employed for 2D projection of 1280D feature vectors in model interpretability analysis. The n_neighbors controls local/global structure balance, min_dist affects cluster tightness, and random_state ensures reproducibility.
    3. 3.
      Implementation: Based on PyTorch 1.12+ framework